# Constrained LTL Specification Learning from Examples

### v1

@@ Abstract
Temporal logic specifications play an important role in a wide range of software analysis tasks, such as model checking, automated synthesis, program comprehension, and runtime monitoring. Given a set of positive and negative examples, specified as traces, LTL learning is the problem of synthesizing a specification, in linear temporal logic (LTL), that evaluates to true over the positive traces and false over the negative ones. In this paper, we propose a new type of LTL learning problem called constrained LTL learning, where the user, in addition to positive and negative examples, is given an option to specify one or more constraints over the properties of the LTL formula to be learned. We demonstrate that the ability to specify these additional constraints significantly increases the range of applications for LTL learning, and also allows efficient generation of LTL formulas that satisfy certain desirable properties (such as minimality). We propose an approach for solving the constrained LTL learning problem through an encoding in a first-order relational logic and reduction to an instance of the maximal satisfiability (MaxSAT) problem. An experimental evaluation demonstrates that ATLAS, an implementation of our proposed approach, is able to solve new types of learning problems while performing better than or competitively with the state-of-the-art tools in LTL learning.
@@ I Introduction
Temporal logic (TL) specifications are a class of specification notations that are used to specify how a system behaves over time. TL specifications, such as linear temporal logic (LTL) [1], have been applied in a wide range of software analysis tasks, such as model checking [2], reactive synthesis [3], program understanding [4], and runtime monitoring [5]. Despite its utility, formalizing and specifying desired system properties in temporal logic is a notoriously challenging and error-prone process [6, 7, 8].
One active area of research that has potential to overcome this challenge is specification learning, where the goal is to automatically infer formal specifications from example traces [4, 9, 8, 6, 10]. These traces may be derived from the actual system executions or manually created by developers as examples. The extracted specification could then be used by developers to comprehend and debug the behavior of a system, document its expected properties, and perform verification through methods such as model checking and theorem proving [6, 11, 12].
@@
For example, consider a mobile navigation robot that moves around three regions, A, B, and R (the last one being a dangerous area that the robot should avoid). Suppose that the robot developer wishes to use a set of formal specifications, to precisely document the robot behavior and to leverage them for verification; an LTL learning tool can be used to extract specifications from traces representing sample behaviors of the robot. Specifically, consider the following two sample traces collected from an operation log: $\langle\textsf{A},\textsf{B},\textsf{B},\ldots\rangle$ (positive; the robot visiting A and then B) and $\langle\textsf{B},\textsf{R},\textsf{R},\ldots\rangle$ (negative; the robot enters R and never leaves, indicating a potentially unsafe behavior). A specification that may be learned from these traces is: $\textsf{A}\land\mathbf{X}(\mathbf{G}\textsf{B})$, i.e., the robot should visit A initially and then stay in B. Technically, this formula is a valid solution to the learning problem. However, it also fails to capture the developer’s intent, i.e., to prohibit the unsafe scenario. Instead, a more ideal specification would constrain the robot from being stuck at R, while not requiring it to start in region A.
Providing an additional set of traces, covering a more diverse set of system behaviors, may be one way to overcome this issue. However, determining a sufficient set of traces to precisely capture the desired characteristics of an LTL specification is challenging. For example, one may wish to extract a specification following a certain syntactic pattern (e.g., a safety or liveness property), minimize the number of times that a certain proposition appears, or learn a formula that is close to another specification (e.g., in the context of specification repair). LTL learning approaches that rely solely on examples lack the expressive power to allow fine-grained control over the characteristics of the formula to be inferred.
@@
The primary target users of ATLAS are expert specifiers and tool builders. For experts in formal logic who currently use LTL learning tools, ATLAS provides them with a more powerful way to control the learning process to extract specifications that precisely capture their intent. Moreover, the generality of FOL constraints offers ATLAS the potential to serve as an underlying, general-purpose engine for other specification-based tools that rely on LTL learning (e.g., specification weakening or repair, as described in Section VII). In addition, a tool built upon ATLAS could automatically convert domain-specific constraints or requirements in natural language to FOL, enabling non-expert users to also benefit from LTL learning.
@@ II Motivating Example
Consider a system analyst working on analyzing a multi-processing algorithm developed by other members of a development team. The analyst plans to employ formal methods but the algorithm is designed without a set of formal specifications. This lack of formal specification is common in practice; even in safety-critical domains, many systems do not have formally defined requirements until the later testing phase, when a need for formal specification to support rigorous verification arises [7]. Although the analyst has a background in formal logic, without an understanding of the details of the algorithm, it would still be challenging to extract a specification that is not only consistent with the algorithm but also representative of its behavior. Specification mining is one approach to overcome this bottleneck—by generating executions (from operation logs or test generation), one can use LTL learning to infer specifications describing the system behavior. These specifications can later be used for documentation, regression testing, or formal verification [7, 6].
@@
The analyst can use an LTL learning tool to infer a specification. When an existing learner such as Flie [9] is applied to the trace in Figure 3, one possible output formula is $a1_{0}$. This formula is technically valid, as it is consistent with this particular execution, but is also arguably not descriptive of the overall algorithm behavior. Providing more traces may not yield a different solution, as every possible trace starts with a state where $a1_{0}$ is true.
One approach for learning more meaningful formulas is to allow the user to constrain the space of candidate formulas to be explored. For instance, suppose that the analyst wants to learn a safety invariant in $\mathbf{G}\phi$ (i.e., $\phi$ always holds). One way to achieve this is to use an existing tool (such as LTLSketcher [8]) that takes a specification template as an additional input. Although templates are a significant improvement, they also fall short of providing fine-grained control that may be desired. For example, without further constraining $\phi$ (e.g., $\phi$ should not contain temporal operators), a template-based learner may return $\mathbf{G}\mathbf{F}(a1_{0})$, which is a valid liveness property but not a safety invariant as intended by the analyst.
@@
Thus, additional constraints and objectives are needed to express the user’s preferences on what a more meaningful safety invariant would be like. For example, consider the following requirement:
“Learn a formula that (1) is in form $\mathbf{G}\phi$ where $\phi$ does not contain temporal operators and (2) maximizes the use of $cs_{0}$ and $cs_{1}$ in $\phi$”.
It first defines a constraint that the formula needs to be a safety invariant. Then, it defines an objective to maximize the use of certain propositions that the analyst is more interested in (i.e., the critical sections in the code). As explained in Section IV, it can be expressed as constraints in FOL, and our approach, ATLAS, can find formula $\mathbf{G}(\neg(cs_{0}\land cs_{1}))$ indicating a mutual exclusion property. However, such constraints and objectives cannot be expressed using existing tools.
@@
As another example, consider the trace shown in Figure 4. This trace is considered negative, as it depicts an undesirable behavior where the execution is stuck at state $(a3_{0},a3_{1})$ (i.e., both processes are waiting for the other to reset the flag). In this case, the analyst wants to learn a liveness property that would rule out this behavior, in the common form of $\mathbf{G}(\phi\rightarrow\mathbf{F}\psi)$. Again, given this pattern as a template, a tool like LTLSketcher [8] may return $\mathbf{G}(a3_{0}\rightarrow\mathbf{F}a1_{0})$. This is a valid liveness property; however, it is also rather unintuitive as it requires a process to enter $a3_{0}$ before $a1_{0}$, even though $a1_{0}$ occurs before $a3_{0}$ in the algorithm’s procedure.
@@
This constraint requires that, for example, if $\phi=a3_{0}$, then $\psi$ should only be either $cs_{0}$ or $a4_{0}$. Given this constraint as an input, ATLAS generates formula $\mathbf{G}(a3_{0}\rightarrow\mathbf{F}cs_{0})$, which is arguably more intuitive than the above one. This formula corresponds to a deadlock-free property.
@@ III-C Alloy and AlloyMax
Alloy [14] is a modeling language based on a first-order relational logic with transitive closure. With its SAT-based engine, the Alloy Analyzer has been applied to a wide range of problems, including protocol verification [20, 21], test generation [22], and bug finding [23, 24].
@@ IV-A Constrained LTL Learning
For the syntactic constraint $\Phi$, a user could specify it to learn an invariant $\mathbf{G}\phi$ where $\phi$ is a propositional formula, a liveness property $\mathbf{G}(\phi\rightarrow\mathbf{F}\psi)$, or a GR(1) formula that is widely used in reactive synthesis [26]. Many of these cannot be expressed by existing tools. Moreover, for the objective $\Psi$, a user could specify it to optimize the syntactic structure in certain way, e.g., minimizing the size of the formula.
@@ IV-C Optimization Objective $\Psi$
The syntax for optimization objective $\Psi$ is an extension to the syntax for the constraint defined in Figure 5:
@@
  • $min(s)$, where $s$ is a non-empty set with a minimal number of elements.
  • $max(s)$, where $s$ is a non-empty set with a maximal number of elements.
@@
An optimization operator can have a superscript $k\in\mathbb{N}^{+}$ (e.g., $min^{k}(s)$) indicating its priority. An objective with a higher priority should be optimized before all the other objectives with lower priorities. When the superscript $k$ is omitted, the constraint has the lowest priority $k=1$.
@@
It minimizes the size of the children relations (edges in a syntax tree), which, in turn, minimizes the size of the formula.
@@ V Technical Approach
The idea of our encoding is inspired by the SAT encoding proposed by Neider and Gavran [9], which is inspired by bounded model checking [27]. The encoding assumes all traces in the sample set are ultimately periodic, known as lasso traces. Such a trace can be represented as $uv^{\omega}$ where $u\in(2^{AP})^{*}$ and $v\in(2^{AP})^{+}$. The observation is that, given a finite set of $AP$, the set of states visited by $uv^{\omega}$ (starting from any time point $t$) is finite and can be determined based only by the finite prefix $uv$. Therefore, this enables us to check LTL temporal operators w.r.t. infinite traces [9]. In addition, the user provides the max number of sub-formulas to bound the search space. Furthermore, we employ several heuristics, described in later sections, to make our $\textnormal{Alloy}^{\textnormal{Max}}$ encoding more succinct and efficient.
@@ V-A Syntax Encoding
We model the syntax of an LTL formula as a directed acyclic graph (DAG). Essentially, a syntax DAG is a syntax tree with shared common sub-formulas. This helps to reduce the search space. The following code snippet shows how this syntax graph can be modeled in $\textnormal{Alloy}^{\textnormal{Max}}$:
@@
Line 2 defines the parent signature for all nodes in a DAG, and line 4 defines the left/right child relations, $l$ and $r$. Line 7 constrains each node to have no paths to itself, i.e., the graph should be acyclic. Specifically, this is achieved by computing all the descendent nodes of $n$ through transitive closure ${}^{\wedge}(l+r)$ as defined in Section IV-B, where $+$ stands for union of sets.
Lines 9-19 define the sub-types of nodes including operators and atomic propositions. Specifically, we add constraints for: binary operators to have exactly one left and one right child (lines 9-12), unary operators to have only left child (line 13-16), and atomic propositions to have no children (line 17-19). Finally, lines 20-21 define a helper function for accessing the root node. Therefore, relation $l$ and $r$ and function $root$ construct the syntax of the learned formula, $\textsf{syntax}(\phi)=\langle\mathcal{L,R},root\rangle$.
@@ V-B Semantics Encoding
This section describes the encoding for the semantics of LTL. We show only the definitions for $\lor,\mathbf{X},\mathbf{F},~{}\text{and}~{}\mathbf{U}$. Encoding for the other operators can be inferred by taking them as reference.
@@
Lines 8-9 define two helper functions seqIndices and futureIdx. While details are omitted, their functionalities are:
@@
For instance, Figure 6 shows an example lasso trace $uv^{\omega}$. The black solid arrows represent relation $next=\{(0,1),(1,2)\}$, and the dashed arrow represents relation $lasso=\{(2,1)\}$ of this trace. Based on the $next$ relation, for time $i=0$ and $i=1$, the next time points are $i^{\prime}=1$ and $i^{\prime}=2$, respectively. For time $i=2$, the state value at $i^{\prime}=3$ equals to the value at $i^{\prime}=1$ because of the repeating $v$’s, and $i^{\prime}=1$ can be retrieved by the $lasso$ relation. Therefore, $\{(2,3),(3,4),\ldots\}$ are unnecessary tuples (shown in Figure 6), and we only need to model the finite prefix $uv$ of a lasso trace.
@@ V-C Problem-Specific Encoding
The following code snippet is the $\textnormal{Alloy}^{\textnormal{Max}}$ template for the problem-specific encoding, particularly the values of all the positive and negative traces.
@@
In this template, the math notations in comments /*..*/ will be replaced by the actual values of a specific learning problem $\langle AP,\mathcal{S},\Phi,\Psi\rangle$. Line 1 defines all the atomic propositions in $AP$ as literal DAG nodes. Lines 2-6 define the finite set of time points given the max length of traces in the sample $\mathcal{S}$ and explicitly specify the initial time $i_{0}$ and the $next$ relation. This leverages a heuristic of Alloy to reduce the search space and improve performance from partial assignment [28].

### v2

@@ Abstract
Temporal logic specifications play an important role in a wide range of software analysis tasks, such as model checking, automated synthesis, program comprehension, and runtime monitoring. Given a set of positive and negative examples, specified as traces, LTL learning is the problem of synthesizing a specification, in linear temporal logic (LTL), that evaluates to true over the positive traces and false over the negative ones. In this paper, we propose a new type of LTL learning problem called constrained LTL learning, where the user, in addition to positive and negative examples, is given an option to specify one or more constraints over the properties of the LTL formula to be learned. We demonstrate that the ability to specify these additional constraints significantly increases the range of applications for LTL learning, and also allows efficient generation of LTL formulas that satisfy certain desirable properties (such as minimality). We propose an approach for solving the constrained LTL learning problem through an encoding in first-order relational logic and reduction to an instance of the maximal satisfiability (MaxSAT) problem. An experimental evaluation demonstrates that ATLAS, an implementation of our proposed approach, is able to solve new types of learning problems while performing better than or competitively with the state-of-the-art tools in LTL learning.
@@ I Introduction
Temporal logic (TL) specifications are a class of specification notations that are used to specify how a system behaves over time. TL specifications, such as linear temporal logic (LTL) [1], have been used in a wide range of software analysis tasks, such as model checking [2], reactive synthesis [3], program understanding [4], and runtime monitoring [5]. Despite its utility, formalizing and specifying desired system properties in temporal logic is a notoriously challenging and error-prone process [6, 7, 8].
One active area of research that has the potential to overcome this challenge is specification learning, where the goal is to automatically infer formal specifications from example traces [4, 9, 8, 6, 10]. These traces may be derived from system executions or manually created by developers as examples. The extracted specification could then be used by developers to comprehend and debug the behavior of a system, document its expected properties, or perform verification through methods such as model checking and theorem proving [6, 11, 12].
@@
For example, consider a mobile navigation robot that moves around three regions, A, B, and R (the last one being a dangerous area that the robot should avoid). Suppose that the robot developer wishes to use formal specifications to precisely document the robot behavior and to leverage them for verification. An LTL learning tool can be used to extract such specifications from traces representing sample behaviors of the robot. Specifically, consider the following two sample traces collected from an operation log: $\langle\textsf{A},\textsf{B},\textsf{B},\ldots\rangle$ (positive; the robot visiting A and then B) and $\langle\textsf{B},\textsf{R},\textsf{R},\ldots\rangle$ (negative; the robot enters R and never leaves, indicating a potentially unsafe behavior). A specification that may be learned from these traces is: $\textsf{A}\land\mathbf{X}(\mathbf{G}\textsf{B})$, i.e., the robot should visit A initially and then stay in B. Technically, this formula is a valid solution to the learning problem. However, it also fails to capture the developer’s intent, i.e., to prohibit the unsafe scenario. Instead, a more ideal specification would constrain the robot from being stuck at R, while not requiring it to start in region A.
Devising an additional set of traces to cover a more diverse set of system behaviors may be one way to overcome this issue. However, in general, determining a set of traces that precisely capture the desired characteristics of a specification is a challenging task. For example, one may wish to extract a specification that fits into a certain pattern or a class (e.g., a safety or liveness property), minimize the number of times that a certain proposition appears, or learn a formula that is close to another specification (e.g., in the context of specification repair). LTL learning approaches that rely solely on examples lack the expressive power to allow fine-grained control over the characteristics of the formula to be inferred.
@@
The primary target users of ATLAS are expert specifiers and tool builders. For experts in formal logic who currently use LTL learning tools, ATLAS provides them with a more powerful way to control the learning process to extract specifications that precisely capture their intent. Moreover, the generality of FOL means that ATLAS has the potential to serve as an underlying, general-purpose engine for other specification-based tools that rely on LTL learning (e.g., specification weakening or repair, as described in Section VII). In addition, a tool built upon ATLAS could automatically convert domain-specific constraints or requirements in natural language to FOL, enabling non-expert users to also benefit from LTL learning.
@@ II Motivating Example
Consider a system analyst working on analyzing a multi-processing algorithm, developed by other engineers. The analyst plans to employ formal verification but the algorithm is designed without a formal specification. This lack of specification is common in practice; even in safety-critical domains, many systems do not have formally defined requirements until a later testing phase, when the need to support rigorous verification arises [7]. Although the analyst has a background in formal logic, without an in-depth understanding of the algorithm, it would be challenging to extract a specification that is not only consistent with the algorithm but also representative of its behavior. Specification mining is one approach to overcome this bottleneck—by generating executions (e.g., from an operation log), one can use LTL learning to infer specifications that capture the system behavior. These specifications can then be used for documentation, regression testing, or formal verification [7, 6].
@@
The analyst may use an LTL learning tool to infer a specification from traces. When an existing learner such as Flie [9] is applied to the trace in Figure 3, one possible output formula is $a1_{0}$. This formula is technically valid, as it is consistent with this particular execution, but is also arguably not descriptive of the overall algorithm behavior. Providing more traces may not yield a different solution, as every possible trace starts with a state where $a1_{0}$ is true.
One approach for learning more meaningful formulas is to allow the user to constrain the space of candidate formulas to be explored. For instance, suppose that the analyst wants to learn a safety invariant of form $\mathbf{G}\phi$ (i.e., $\phi$ always holds). One way to achieve this goal is to use a learner that takes a specification template as an additional input (such as LTLSketcher [8]). Although templates are a significant improvement, they also fall short of providing fine-grained control that may be desired. For example, without further constraining $\phi$ (e.g., $\phi$ should not contain temporal operators), a template-based learner may return $\mathbf{G}\mathbf{F}(a1_{0})$, which is a valid liveness property but not a safety invariant as intended by the analyst.
@@
We argue that additional constraints and objectives are needed to express the analyst’s preference on what a meaningful safety invariant looks like. For example, consider the following requirement:
“Learn an LTL formula that is in form $\mathbf{G}\phi$ where $\phi$ does not contain temporal operators, and maximizes the use of $cs_{0}$ and $cs_{1}$ in $\phi$”.
It first defines a constraint that the formula needs to be a safety invariant. Then, it defines an objective to maximize the use of certain propositions that the analyst is interested in (i.e., the critical sections in the code). As explained in Section IV, it can be expressed as constraints in FOL, and our approach, ATLAS, can find formula $\mathbf{G}(\neg(cs_{0}\land cs_{1}))$, which captures a mutual exclusion property. Such constraints and objectives cannot be expressed using the existing learning tools.
@@
As another example, consider the trace shown in Figure 4. This trace is considered negative, as it depicts an undesirable behavior where the execution is stuck at state $(a3_{0},a3_{1})$ (i.e., both processes are waiting for the other to reset the flag). In this case, the analyst wants to learn a liveness property that would rule out this behavior, in form of $\mathbf{G}(\phi\rightarrow\mathbf{F}\psi)$. Again, given this pattern as a template, a tool like LTLSketcher [8] may return $\mathbf{G}(a3_{0}\rightarrow\mathbf{F}a1_{0})$. This is a valid liveness property; however, it is also rather unintuitive as it requires a process to enter $a3_{0}$ before $a1_{0}$, even though $a1_{0}$ occurs before $a3_{0}$ in the algorithm’s procedure.
@@
This constraint requires that, for example, if $\phi=a3_{0}$, then $\psi$ should be either $cs_{0}$ or $a4_{0}$. Given this constraint as an input, ATLAS generates formula $\mathbf{G}(a3_{0}\rightarrow\mathbf{F}cs_{0})$, which is arguably more intuitive than the above one. This formula corresponds to a deadlock-free property.
@@ III-C Alloy and AlloyMax
Alloy [14] is a modeling language based on first-order relational logic with transitive closure. With its SAT-based engine, the Alloy Analyzer has been applied to a wide range of problems, including protocol verification [20, 21], test generation [22], and bug finding [23, 24].
@@ IV-A Constrained LTL Learning
For the syntactic constraint $\Phi$, a user could specify it to learn an invariant $\mathbf{G}\phi$ where $\phi$ is a propositional formula, a liveness property $\mathbf{G}(\phi\rightarrow\mathbf{F}\psi)$, or a GR(1) formula, which is widely used in reactive synthesis [26]. Many of these cannot be expressed by existing tools. Moreover, for the objective $\Psi$, a user could specify it to optimize the syntactic structure in certain way, e.g., minimizing the size of the formula.
@@ IV-C Optimization Objective $\Psi$
The syntax for optimization objective $\Psi$ is an extension to the syntax defined in Figure 5:
@@
  • $min(s)$, where $s$ is a non-empty set and the number of elements in $s$ should be minimized.
  • $max(s)$, where $s$ is a non-empty set and the number of elements in $s$ should be maximized.
@@
An optimization operator can be assigned a superscript $k\in\mathbb{N}^{+}$ (e.g., $min^{k}(s)$) indicating its priority. An objective with a higher priority should be optimized before all the other objectives with lower priorities. When the superscript $k$ is omitted, the constraint has the lowest priority $k=1$.
@@
It minimizes the size of the child relations (edges in a syntax tree), which, in turn, minimizes the size of the formula.
@@ V Technical Approach
The idea of our encoding is inspired by the SAT encoding proposed by Neider and Gavran [9], which is inspired by bounded model checking [27]. The encoding assumes that all traces in the sample set are ultimately periodic, known as lasso traces. Such a trace can be represented as $uv^{\omega}$ where $u\in(2^{AP})^{*}$ and $v\in(2^{AP})^{+}$. The observation is that, given a finite set of $AP$, the set of states visited by $uv^{\omega}$ (starting from any time point $t$) is finite and can be determined based only on the finite prefix $uv$. Therefore, this enables us to check LTL temporal operators w.r.t. infinite traces [9]. In addition, the user provides the maximum number of sub-formulas to bound the search space. Furthermore, we employ several heuristics, described in later sections, to make our $\textnormal{Alloy}^{\textnormal{Max}}$ encoding more succinct and efficient.
@@ V-A Syntax Encoding
We model the syntax of an LTL formula as a directed acyclic graph (DAG). Essentially, a syntax DAG is a syntax tree with shared common sub-formulas. This helps reduce the search space. The following code snippet shows how this syntax graph can be modeled in $\textnormal{Alloy}^{\textnormal{Max}}$:
@@
Line 2 defines the parent signature for all nodes in a DAG, and line 4 defines the left/right child relations, $l$ and $r$. Line 7 constrains each node to have no paths to itself, i.e., the graph should be acyclic. Specifically, this is achieved by computing all the descendent nodes of $n$ through transitive closure ${}^{\wedge}(l+r)$ as defined in Section IV-B, where $+$ is the union operator.
Lines 9-19 define the sub-types of nodes including operators and atomic propositions. Specifically, we add constraints for: binary operators to have exactly one left and one right child (lines 9-12), unary operators to have only left child (line 13-16), and atomic propositions to have no children (line 17-19). Finally, lines 20-21 define a helper function for accessing the root node. Therefore, relation $l$ and $r$, and function $root$ construct the syntax of the learned formula, $\textsf{syntax}(\phi)=\langle\mathcal{L,R},root\rangle$.
@@ V-B Semantics Encoding
This section describes the encoding for the semantics of LTL. We show only the definitions for $\lor,\mathbf{X},\mathbf{F},~{}\text{and}~{}\mathbf{U}$. Encodings for the other operators can be defined in a similar manner.
@@
Lines 8-9 define two helper functions seqIndices and futureIdx. Their functionalities are:
@@
For instance, Figure 6 shows an example lasso trace $uv^{\omega}$. The solid black arrows represent relation $next=\{(0,1),(1,2)\}$, and the dashed arrow represents relation $lasso=\{(2,1)\}$ of this trace. Based on the $next$ relation, for time $i=0$ and $i=1$, the next time points are $i^{\prime}=1$ and $i^{\prime}=2$, respectively. For time $i=2$, the state value at $i^{\prime}=3$ equals to the value at $i^{\prime}=1$ because of the repeating $v$’s, and $i^{\prime}=1$ can be retrieved by the $lasso$ relation. Therefore, $\{(2,3),(3,4),\ldots\}$ are unnecessary tuples (shown in Figure 6), and we only need to model the finite prefix $uv$ of a lasso trace.
@@ V-C Problem-Specific Encoding
The following code snippet shows the $\textnormal{Alloy}^{\textnormal{Max}}$ template for the problem-specific encoding, particularly the values of all the positive and negative traces.
@@
In this template, the mathematical expressions in the comments /*..*/ will be replaced by the actual parameters of a specific learning problem $\langle AP,\mathcal{S},\Phi,\Psi\rangle$. Line 1 defines all the atomic propositions in $AP$ as literal DAG nodes. Lines 2-6 define the finite set of time points given the maximum length of traces in the sample $\mathcal{S}$ and explicitly specify the initial time $i_{0}$ and the $next$ relation. This leverages a heuristic in Alloy to improve performance through partial instances [28].